OpenStudy (anonymous):

4[2ln(x-2)+(1/2)ln(x+1)-ln(x^2-8)] Needs to be written in condensed form

OpenStudy (anonymous):

ln (x) + ln (y) - ln (z) = ln (xy/z) a ln(x) =ln (x)^a